OpenStudy (anonymous):

Which expression is equivalent to b^m * b^n? A) b^m+n b) b^m-n c) b^m*n D) b^m/n

OpenStudy (freckles):

if you have a m amount of factors of b and you put in n amount of factors of b then how many factors of b do you have?

OpenStudy (freckles):

let's do an example what is 2^3 * 2^4 ?

OpenStudy (freckles):

do not evaluate write as 2^some number

OpenStudy (anonymous):

48?

OpenStudy (freckles):

2^3 means you have 3 factors of 2 2^4 means you have 4 factors of 2 when you multiply things together, the things you are multiplying are called factors if you do 2^3 * 2^4 you are saying you are multiplying 3 factors of 2 to 4 factors of 2 how many factors of 2 does this mean you have?

DivineSolar (divinesolar):

Evaluate for b=2,m=3,n=4 (2^3)(2^4) (2^3)(2^4) =128

DivineSolar (divinesolar):

Thats for b^m * b^n with b=2, m=3, n=4

DivineSolar (divinesolar):

Evaluate for b=2,m=3,n=4 2^3+4 2^3+4 =12

DivineSolar (divinesolar):

Is that the same?

OpenStudy (agent0smith):

@divinesolar you misinterpreted the options. They're all exponents, the asker just left out parentheses